# SODAs: Sparse Optimization for the Discovery of Differential and Algebraic Equations

### v1

@@ 4.1 Example 1: Chemical Reaction Networks
    In each application of SODAs, Savitzky-Golay filters were again applied [48] to smooth noisy derivatives. For CRN2, Pareto front analysis correctly suggested three refinements, while, for CRN3, it correctly suggested five (see supplementary Section 2.3). We again assessed the response of the method to increasingly noisy data across a variety of resolutions at fixed polynomial degree $p=2$ (Figure 3c). While we find a similar scaling of the data requirement with respect to noise as in Figure 3b, there is a notable difference in how the data requirement responds to an increase in underlying model complexity vs. an increase in library degree, even if the two modifications lead to $\Theta$ of similar cardinality. For example, $p=2$ recovery for CRN2 requires significantly more data points per initial condition than $p=3$ recovery for CRN1 at $>8\%$ noise, despite the two cases having $|\Theta|=35$ and $|\Theta|=34$, respectively. From this we may hypothesize that SODAs performance is more sensitive to the library degree than the dimension of the library’s monomial basis, a notion consistent with our earlier discussions on the impact of higher-order terms on library correlation. Our results also showed that it is not necessarily the case that more complex underlying models always have higher data requirements for rediscovery, as there were regimes where CRN2 required fewer data points per initial condition than CRN1, and regimes where CRN3 required less data than CRN2 despite both requiring more than CRN1. This points towards the importance of data sampling; simply the quantity of the data alone does not determine its ability to inform sparse-model selection, and, for example, the types of dynamics sampled within the data may be important to model discovery.
@@ 4.2 Example 2: Non-linear and Chaotic Pendulums
    In this example, we demonstrate the capability of SODAs to identify a reduced coordinate system from pixel data obtained from video footage of two different types of pendulums across three different scenarios: 1) real-time footage of a single pendulum exhibiting non-linear behavior through large-angle swings, 2) video recordings of the animation of a highly damped pendulum, and 3) video recordings of the animation of a double pendulum exhibiting chaotic behavior. Screen recordings of the animation were used for 2 and 3 to mimic the process of extracting data from video, even though a simulation was used to produce the animation. These cases illustrate the use of SODAs in mechanical systems, where state measurements are obtained through images, videos, or sensor data.
    When modeling the motion of mechanical systems, there often exist coordinate transformations that lead to simple equations explaining the motion. For example, the natural coordinate system for modeling the motion of a single pendulum is known to be the one-dimensional polar coordinate system, i.e., to model the angle $\omega$ that the pendulum makes with a reference axis in the 2D plane (Figure 4a - top) or in the case of a double pendulum, the two angles, $\omega_{1},\omega_{2}$ , that the two pendulums make with the reference axis (Figure 4a - bottom). However, finding the appropriate coordinate transformation to produce the most parsimonious system is often challenging and not clear from raw measurements of the states. Recently, several data-driven methods have been used to find reduced coordinates, for example, using neural networks to learn nonlinear mappings to a lower-dimensional state-space [62, 63]. Here, we show how SODAs can assist in this process using the algebraic finder step. Specifically, discovering the algebraic relationships between state variables can give intuition about the nature of symmetries within the system and the natural coordinate system suitable for obtaining a simple model.
    In all three cases presented here, SODAs identify the intrinsic algebraic constraints in the system, implying the reduced polar coordinate system inherent to it, solely based on pixel data that represents the time series of the pendulum’s position in Cartesian coordinates $(x,y)$. For Cases 1 and 2, once the coordinate transformation is identified, the equation of the pendulum is discovered in the new transformed coordinate system.
@@
    where $m$ and $l$ represent the mass and length of the pendulum respectively, $\omega$ denotes the angle that the pendulum makes with the vertical axis, $g$ represents the acceleration due to gravity, and $\alpha$ is the damping coefficient which takes the value of zero when damping is negligible. The physical constraint of the single pendulum can be captured by the equation
@@
    A double pendulum consists of two pendulums attached end to end, as shown in Figure 4a - botom. This assembly of two pendulums is well-studied and known to exhibit chaotic motion [64]. To simplify the analysis, we will neglect damping. The equations of motion for a double pendulum, derived using Lagrangian mechanics [64], are given by:
@@
    Figure 4: Example 3: Application to non-linear pendulum. (a) Schematic diagram: single pendulum (top) and double pendulum (bottom). (b) Scaled pixel data: single pendulum experiments (left) and double pendulum animation (right). (c) Data requirement: comparison between damped single pendulum and double pendulum. (d) SVD analysis: determining the number of algebraic constraints in the case of the chaotic double pendulum with degree 5 library.
@@
    For this study, pixel data from three different cases were used: 1) single pendulum with negligible damping: video recording of single pendulum experiments with negligible damping and three different initial conditions (Figure 4b - left), 2) heavily damped pendulum: video recording of animation of a heavily damped pendulum with four different initial conditions, 3) chaotic double pendulum: single video recording of the animation of a double pendulum exhibiting chaotic behavior (Figure 4b - right). Details about the collection of the pixel data and preprocessing can be found in supplementary Section 1.2.
@@
    SODAs successfully discovered the constraint (29) for Cases 1 and 2. Note that $l=1$ for our dataset due to the scaling of the pixel data. After discovering the constraint and identifying the reduced polar coordinate system, we applied the following transformation to move to the polar coordinate system:
@@
    where $\phi$ is the angle formed by the mass of the pendulum with the positive $x$axis (Figure 4a - top). We made this choice instead of the natural one in(30) because, given (29), transformation (35) is the most straightforward coordinate transformation in the absence of additional information about the system. Indeed, under the translation $\phi=\omega+\frac{3\pi}{2}$, it can be easily verified that the equations we discovered after applying the transformation (35) are consistent with (30) and (28). To discover the dynamic equations in the polar coordinate system, we applied a Savitzky-Golay filter [48] with a window size of 12 to compute the first and second derivatives of the $\phi$. The candidate library consisted of monomials of $\phi,\dot{\phi},\cos\phi$ and $\sin\phi$ : $\Theta_{*}=\{\phi,\dot{\phi},\cos\phi,\sin\phi,\phi^{2},\phi\dot{\phi},\ldots\}$. Using (35), we included $x$ and $y$ in the library instead of calculating $\cos\phi$ and $\sin\phi$ directly from $\phi$. Similar to the algebraic finder step, we restricted the monomial degree to 3 for Case 1 and varied it between $2$ and $5$ for Case 2 for a comparative study. After using sparse regression, where sequential threshold with inner iterations consisting of LASSO regularization was applied, we used a curve fit package to determine the correct parameters. In both case 1 and case 2, simulated solutions based on the discovered equations were able to predict the test pixel data accurately (see supplementary Section 3 for details on discovered equations and parameter estimation).
@@
    For Case 3, SODAs robustly identified the constraints (33)–(34) from pixel data obtained from a single video segment, using various upper bounds on the degree of the monomial library. Note that the polar nature of the states $(x_{2},y_{2})$ is not immediately evident from the pixel data of the secondary pendulum in the double-pendulum (Figure 4b - right). This is because the secondary pendulum does not have a fixed pivot and sweeps trajectory in the 2D plane that doesn’t correspond to an identifiable pattern. Nevertheless, SODAs discovered an expanded version of (34). This algebraic discovery enables us to determine the transformation to the polar coordinate system: $\phi_{1}=\arctan\frac{y_{1}}{x_{1}}$ and $\phi_{2}=\arctan\frac{y_{1}-y_{2}}{x_{1}-x_{2}}$, which can be further used for dynamic discovery. The dynamic discovery of equations governing the motion of the double pendulum similar to the form (31) –(32) based on polar coordinates, has been successfully demonstrated using SINDy-based methods [27]. We do not repeat this work here using our pixel data because it requires addressing noisy derivatives and may require advanced smoothing filters to include higher-order derivatives of $\phi_{1}$ and $\phi_{2}$ in the candidate library. Instead, we use this example to demonstrate how SVD analysis, as discussed in Section 3, can be used to track the number of algebraic relationships present in the library and decide when to stop the library refinement in the algebraic finder step.

### v2

@@ 4.1 Example 1: Chemical Reaction Networks
Testing recovery in the more complex CRN2 and CRN3 examples demonstrates how SODAs responds to a larger state space, richer dynamics, and, as a consequence, more complex algebraic relationships (Figure 3). Savitzky-Golay filters were again applied [Bromba and Ziegler [15]] to smooth noisy derivatives. For CRN2, Pareto front analysis correctly suggested three refinements, while, for CRN3, it correctly suggested five (see Supplementary-section S3(e)). We assessed the method for increasingly noisy data across a variety of resolutions at fixed polynomial degree $p=2$ (Figure 3c). The data requirement increases with noise, but there is a notable difference in how the data requirement responds to an increase in underlying model complexity vs. an increase in library degree, even if the two modifications lead to $\Theta$ of similar cardinality. For example, $p=2$ recovery for CRN2 requires significantly more data points per initial condition than $p=3$ recovery for CRN1 at $8\%$ noise, despite the two cases having $|\Theta|=35$ and $|\Theta|=34$, respectively. We hypothesize that SODAs performance is more sensitive to the library degree than the dimension of the library’s monomial basis, a notion consistent with our earlier discussions on the impact of higher-order terms on library correlation. Our results also showed that it is not necessarily the case that more complex underlying models always have higher data requirements for rediscovery, as there were regimes where CRN2 required fewer data points per initial condition than CRN1, and regimes where CRN3 required fewer data points than CRN2. This highlights that the quantity of the data alone does not determine success, and the range of dynamics sampled within the data may be important to model discovery.
We finally note that, due to the instability that can arise during forward simulation of discovered DAEs and the inherent difficulty of parameter estimation for DAE systems, we relied solely on the goodness of fit measured by the $R^{2}$ score as the primary metric for model selection in our numerical experiments. This choice enables a consistent and robust comparison of candidate models without requiring successful forward simulation or stable parameter estimation. In practice, however, a practitioner may instead employ a DAE solver together with a parameter estimation framework tailored to the specific class of DAEs under consideration, using simulation-based validation as an additional criterion for model selection among candidates produced by SODAs under different data regimes, noise realizations, library choices, and hyperparameter settings. We provide a concrete demonstration of this approach in Supplementary-section S3(f), where three representative models discovered at $10\%$ additive noise using a degree-2 library are examined: one with the correct structure and two with incorrect structures of varying similarity to the true model. In this example, parameters are estimated, and the resulting DAEs are simulated and compared against noisy data, illustrating that when the discovered model structure deviates substantially from the true system, the simulated DAEs tend to become unstable, and meaningful parameter inference and simulation become difficult.
4.2 Example 2: Power-grid Dynamics
The operation of modern power grids relies on detailed modeling of dynamical behavior and network topology, which are critical for grid stability, efficiency, and control against disturbances [Gong et al. [27]]. Such model information directly supports daily operations, scheduling, and advanced techniques like state estimation, fault monitoring, and optimal power flow computation. However, the rise of renewable generation has led to decentralized power generation, making detailed grid models less accessible in distribution systems In these settings, the power-grid network changes frequently due to the addition or removal of generating units (e.g., rooftop solar, battery storage units) and network components (e.g., lines, switches, transformers), as well as uncertainty in distributed energy resources, line maintenance, faults, and other unplanned events. Yet, traditional power-grid monitoring and fault detection methods are often grounded on parameter estimation techniques, requiring extensive prior knowledge of grid configurations and physical properties [Pierre et al. [57], Gopakumar et al. [28], Liu et al. [38], Gaskin et al. [25]]. The increasing deployment of phasor measurement units (PMUs) [Rodrigues et al. [61]], capable of providing high-resolution data, offers a significant opportunity for automated model discovery [Ardakanian et al. [3], Yuan et al. [83]]. Here, we show that the sparse discovery of differential-algebraic models can robustly infer the network topology, which is encoded by the algebraic relationships among power injections at each bus. Using numerical results, we demonstrate that the proposed method accurately infers network models across a wide range of scenarios, including both small- and large-scale power grids, varying levels of state perturbation, and diverse PMU signal qualities.
System description.
The dynamics of power grids can be modeled as a network of coupled oscillators (nodes) interconnected through power transmission lines (links) [Sadamoto et al. [63], Dörfler et al. [21], Nishikawa and Motter [54]]. The nodes correspond to conventional generators and consumers (geographically represented by substations and their attached load). In this representation, the dynamics of generators are governed by the swing equation [Saadat [62]]:
north
SODAs application.
We considered three IEEE benchmark cases with varying complexity (Figure 4a): IEEE-4, IEEE-9, and IEEE-39. Time series were generated under two distinct conditions: i) strong perturbations (Figure 4b - top), which excite nonlinear transient behaviors, and ii) weak perturbations (Figure 4b - bottom), suitable for small-signal stability analysis. We assume that the power grid is equipped with phasor measurement units (PMUs) placed on all $N$ buses, providing us with high-resolution, high-frequency data for phase angle $\phi_{i}(t)$, frequency $\dot{\phi}_{i}(t)$, line voltage $V_{i}(t)$, and electrical power $P_{\rm e,i}(t)$ and $Q_{\rm e,i}(t)$. More information on network details, simulations, measurement frequency, and practicality of assumptions can be found in the Supplementary-section S4(a). For the algebraic finder step, STOLS was utilized instead of LASSO followed by thresholding, as the latter indicated a poor fit when initially applied.
SODAs was applied to the simulated time series to identify the network structure and parameters of the governing DAE system as described in equations (23)–(25). By using the well-established physical principles of power-grid dynamics, we achieved a significant advantage in constraining SODAs within a more informed framework. More specifically, the conservation of active power at each node leads to inherent algebraic constraints that we exploited to improve the identification process. In constructing the candidate library, we included the following functional forms: $P_{{\rm e},i}$ (active power states), $\phi_{i}$ (phase angles), $\dot{\phi}_{i}$(frequency), and sine terms of phase differences $\sin(\phi_{i}-\phi_{j})$ for all pairs $(i,j)$. In a network with $N$ nodes, this initial library included $3N+\frac{N(N+1)}{2}$ terms, which amounts to 39, 114, and 1,372 terms for the IEEE-4, IEEE-9, and IEEE-39 benchmarks, respectively. Additionally, by incorporating prior knowledge, we further reduced the number of candidate terms used to discover each conservation law. For example, the conservation law involving $P_{{\rm e},i}$ at a given node $i$ restricts the algebraic relations to include only sine terms of the form $\sin(\phi_{i}-\phi_{j})$, for all nodes $j$. This restriction decreases the number of coupling terms for each node from $\frac{N(N+1)}{2}$ to N1N-1. Consequently, the number of library terms used to discover each power-flow equation consisted of 23, 47, and 195 terms for the IEEE-4, IEEE-9, and IEEE-39 benchmark cases, respectively. Using the Daefiender.FeatureCouplingTransformer class, we efficiently enforced these additional library constraints, minimizing computational complexity and enhancing the accuracy of identified models by eliminating irrelevant terms.
In the algebraic finder step, conservation laws constraining the active powers $P_{{\rm e},i}$ for all $i$ are discovered (25). These algebraic equations inherently reveal the power grid’s network structure (encoded by the admittance matrix $Y$). In the dynamic finder step of SODAs, the differential equations governing the generators are discovered using a SINDy framework. We approximated the left-hand side of the dynamic equations (23)–(24) by utilizing derivatives of spline-fitted phase angles and applied sparse regression techniques, where LASSO was followed by sequential thresholding. Parameter estimation was subsequently performed through linear fitting with the discovered model structure. Note that for the identification of the dynamic equation, we modified the candidate library by removing all $\sin(\phi_{i}-\phi_{j})$ terms, which had already been accounted for by the algebraic constraints. Instead, we incorporated state measurements of $P_{{\rm e},i}$ and hence reducing the number of candidate functions in the dynamic finder step to $3N$. This refinement has also helped to avoid using the coupling terms $\sin(\phi_{i}-\phi_{j})$, which have high sensitivity to noise in the data, for the dynamic finder step. We observed complete recovery of the true ODE system whenever the algebraic constraints were discovered with 100% accuracy. In contrast, attempts to identify coupled ODE equations without segregating the algebraic terms led to poor model recovery, even at SNR as high as 40dB, emphasizing the importance of this methodological separation of algebraic and differential equations.
We present SODAs’ performance on identifying the ground-truth model for all three test cases, with a primary focus on the IEEE-39 benchmark test case (Figure 4c and 4d). We conducted similar analyses on the IEEE-4 and IEEE-9 benchmark networks, where SODAs demonstrated performance comparable or superior to that of the more complex IEEE-39 case. To evaluate the robustness of SODAs, we introduced varying amounts of noise into the simulated time-series data and examined performance across four different signal-to-noise ratios (SNR): no noise, 40 dB, 30 dB, and 20 dB. For each SNR level, we generated five distinct noise realizations and analyzed the performance as an average across these realizations. We found no significant variations in performance among the different noise realizations within the same SNR level, which allowed us to exclude error bars related to noise realizations in our figures.
For SNRs of no noise, 40 dB, and 30 dB, SODAs achieved 100% recovery of all algebraic relationships, as well as the dynamic ODEs, using the refined library for both small and large perturbations. Even under high noise conditions (SNR of 20 dB), SODAs maintained a recovery rate exceeding 80% for algebraic relations with 100 perturbations. The algebraic relationships related to power conservation at generator nodes were identified with relatively small number of perturbations, underscoring the significant role of these nodes in influencing the overall dynamics of the power grid. Techniques such as signal averaging and advanced filtering can improve the SNR, even when the inherent measurement SNR is lower. Therefore, an SNR of 30 dB is assumed to be realistic for our further analysis, as other studies have reported PMU devices with signal quality at or above this level [Brown et al. [16], Bhandari et al. [9], Frigo et al. [24], Zhang et al. [84]]. We evaluated the performance of SODAs at a 30 dB SNR across various benchmark cases and levels of perturbations (Figure 4d). We did a comparative analysis of the three benchmark cases, in terms of the number of perturbations required to recover all true DAE system components, and found that the number of perturbations required for full recovery increased with the complexity of the power grid, yet it remained within the realistic number of permutations observed in practice (Figure 4d - left). Also, strong perturbations that induce significant transients in system dynamics enabled quicker recovery of the DAE system as expected (Figure 4d - right).
Additionally, we also investigated whether the models discovered by SODAs produce physically meaningful power system trajectories when simulated. Across multiple test cases, we found that accurately recovering the algebraic (power-flow) constraints was crucial for generating valid phase angle dynamics. When these algebraic relationships were misidentified, the resulting DAE models often failed to demonstrate frequency synchronization. This failure was evident as the phase angles did not converge to a steady equilibrium in the co-rotating frame, indicating that the predicted system dynamics were qualitatively incorrect. A representative example from the IEEE-4 benchmark system, illustrating the differences between correctly and incorrectly identified algebraic constraints, can be found in Supplementary-section S4(b).
Figure 4: Example 2: Discovery of power-grid networks. (a) Power-grid networks of the IEEE-4 (top left), IEEE-9 (bottom left), and IEEE-39 (right) benchmark systems. Generator buses are represented by blue nodes on the outer circle, while aggregated load buses are represented by gray nodes. (b) Time series of phase dynamics $\phi_{i}(t)$ for all nodes in the IEEE-39 power grid under different levels of perturbations: large perturbations (top), large perturbations zoomed-in (middle), and small perturbations zoomed-in (bottom). (c) Performance comparison of SODAs applied to IEEE-39 benchmark system across different SNR levels and : small perturbations (left) and large perturbations (right). (d) Performance comparison of SODAs at 30 dB SNR across: different benchmark cases (left) and different perturbation strengths (right).
@@ 4.3 Example 3: Non-linear and Chaotic Pendulums
In this example, we demonstrate how the algebraic finder step of SODAs can identify a reduced coordinate system, specifically the polar coordinate system that arises from the physical constraints of three different pendulums. This example differs from the previous two examples, where the goal was to derive a complete set of DAEs that govern the system’s dynamics. In this example, we apply SODAs to analyze pixel data obtained from video recordings of three different types of pendulums: 1) real-time footage of a single pendulum exhibiting non-linear behavior through large-angle swings, 2) video recordings of the animation of a highly damped pendulum, and 3) video recordings of the animation of a double pendulum exhibiting chaotic behavior. Screen recordings of the animation were used for cases 2 and 3 to mimic the process of extracting data from video, even though a simulation was used to produce the animation. These cases illustrate the use of SODAs in mechanical systems, where state measurements are obtained through images, videos, or sensor data.
When modeling the motion of mechanical systems, there often exist coordinate transformations that lead to simple equations explaining the motion. For example, the natural coordinate system for modeling the motion of a single pendulum is known to be the one-dimensional polar coordinate system, i.e., to model the angle $\omega$ that the pendulum makes with a reference axis in the 2D plane (Figure 5a - top) or in the case of a double pendulum, the two angles, $\omega_{1},\omega_{2}$ , that the two pendulums make with the reference axis (Figure 5a - bottom). However, finding the appropriate coordinate transformation to produce the most parsimonious system is often challenging and not clear from raw measurements of the states. Recently, several data-driven methods have been used to find reduced coordinates, for example, using neural networks to learn nonlinear mappings to a lower-dimensional state-space [Mars Gao and Nathan Kutz [43], Yang et al. [81]]. Here, we show how SODAs can assist in this process using the algebraic finder step. Specifically, discovering the algebraic relationships between state variables can give intuition about the nature of symmetries within the system and the natural coordinate system suitable for obtaining a simple model.
In all three cases presented here, SODAs identify the intrinsic algebraic constraints in the system, implying the reduced polar coordinate system inherent to it, solely based on pixel data that represents the time series of the pendulum’s position in Cartesian coordinates $(x,y)$. For Cases 1 and 2, once the coordinate transformation is identified, the ODE equation of the pendulum is discovered in the new transformed coordinate system. For Case 3, we do not recover the dynamic equations. Together, these cases highlight that discovering algebraic constraints can aid in finding a coordinate system and is less data-intensive than discovering the complete DAE system.
@@
where $m$ and $l$ represent the mass and length of the pendulum, respectively, $\omega$ denotes the angle that the pendulum makes with the vertical axis, $g$ represents the acceleration due to gravity, and $\alpha$ is the damping coefficient, which is zero when damping is negligible. The physical constraint of the single pendulum can be captured by the equation
@@
A double pendulum consists of two pendulums attached end to end, as shown in Figure 5a - botom. This assembly of two pendulums is well-studied and known to exhibit chaotic motion [Shinbrot et al. [69]]. To simplify the analysis, we will neglect damping. The equations of motion for a double pendulum, derived using Lagrangian mechanics [Shinbrot et al. [69]], are given by:
@@
Figure 5: Example 3: Application to non-linear pendulum. (a) Schematic diagram: single pendulum (top) and double pendulum (bottom). (b) Scaled pixel data: single pendulum experiments (left) and double pendulum animation (right). (c) Data requirement: comparison between damped single pendulum and double pendulum. (d) SVD analysis: determining the number of algebraic constraints in the case of the chaotic double pendulum with degree 5 library. Y-axes are given in natural log scale.
@@
For this study, pixel data from three different cases were used: 1) single pendulum with negligible damping: video recording of single pendulum experiments with negligible damping and three different initial conditions (Figure 5b - left), 2) heavily damped pendulum: video recording of animation of a heavily damped pendulum with four different initial conditions, 3) chaotic double pendulum: single video recording of the animation of a double pendulum exhibiting chaotic behavior (Figure 5b - right). OpenCV [Bradski and Kaehler [13]] package in Python was used to extract the pixel data from the video recordings. We would like to note that converting raw video or animation frames into reliable image-plane measurements is a complex problem. This process is managed using a standard, well-established computer vision routine implemented with OpenCV’s standard functions for color segmentation, morphology, and contour detection, [Bradski and Kaehler [13], Pulli et al. [58]]. Managing factors such as camera perspective, lens distortion, focus, and noise associated with the video footage is a nontrivial challenge, and hence we relied on these established methods during this stage. More details about the collection of the pixel data and preprocessing can be found in Supplementary-section S5(a).
@@
SODAs successfully discovered the constraint (27) for the single pendulum Cases 1 and 2. Note that $l=1$ for our dataset due to the scaling of the pixel data. After discovering the constraint and identifying the reduced polar coordinate system, we applied the following transformation to move to the polar coordinate system:
@@
where $\phi$ is the angle formed by the mass of the pendulum with the positive $x$ -axis (Figure 5a - top). We made this choice instead of the natural one in(28) because, given (27), transformation (33) is the most straightforward coordinate transformation in the absence of additional information about the system. Indeed, under the translation $\phi=\omega+\frac{3\pi}{2}$, it can be easily verified that the equations we discovered after applying the transformation (33) are consistent with (28) and (26). For this simple case, we have sufficient data to discover the dynamic equations in the polar coordinate system. We applied a Savitzky-Golay filter [Bromba and Ziegler [15]] with a window size of 12 to compute the first and second derivatives of the $\phi$. The candidate library consisted of monomials of $\phi,\dot{\phi},\cos\phi$ and $\sin\phi$ : $\Theta_{*}=\{\phi,\dot{\phi},\cos\phi,\sin\phi,\phi^{2},\phi\dot{\phi},\ldots\}$. Using (33), we included $x$ and $y$ in the library instead of calculating $\cos\phi$ and $\sin\phi$ directly from $\phi$. Similar to the algebraic finder step, we restricted the monomial degree to 3 for Case 1 and varied it between $2$ and $5$ for Case 2 for a comparative study. After ODE model structures were discovered using sequential LASSO and thresholding, we used a curve fit package to determine the correct parameters. In both case 1 and case 2, simulations of the discovered equations predicted the test pixel data accurately (see Supplementary-section S5(b) for details on discovered equations and parameter estimation). Noteably, the discovery of the dynamic equations in the library with $\sin\phi$ and $\cos\phi$ required more data than the algebraic equations in the original Cartesian coordinates for monomial degree 5 (Fig. 5c), likely due to a combination of higher correlations in the candidate library and the need to calculate derivatives from noisy data for ODE discovery.
@@
For Case 3, SODAs robustly identified the constraints (31)–(32) from pixel data obtained from a single video segment, using various upper bounds on the degree of the monomial library. Note that the polar nature of the states $(x_{2},y_{2})$ is not immediately evident from the pixel data of the secondary pendulum in the double-pendulum (Figure 5b - right). Nevertheless, SODAs discovered an expanded version of (32). This algebraic discovery enables us to determine the transformation to the polar coordinate system: $\phi_{1}=\arctan\frac{y_{1}}{x_{1}}$ and $\phi_{2}=\arctan\frac{y_{1}-y_{2}}{x_{1}-x_{2}}$, which could, in principle, be further used for dynamic discovery. The dynamic discovery of equations governing the motion of the double pendulum similar to the form (29) –(30) based on polar coordinates, has been successfully demonstrated using SINDy-based methods [Kaheman et al. [33]]. However, discovering the algebraic constraint from a particular data set does not guarantee the successful recovery of the ODE system in the new coordinate system, and previously successful SINDy-based methods required more data than used here. Success depends on several factors, including the information content in the data relative to the dynamics of the system, noise-robust estimation of derivatives, and more. We suspect discovering this ODE requires weak approaches to handle the higher-order derivatives of $\phi_{1}$ and $\phi_{2}$. Without such methods, we are unable to rediscover the ODE system defined by equations (29)–(30) as part of this example, but note that even when data is insufficient to discover the dynamics, it can still be used to learn about constraints. To demonstrate the robustness of algebraic equation discovery, we show that the SVD analysis (Section 3) can be used to evaluate how many library refinements lead to meaningful equations.